how many feet of boarder will Mrs.Kohl need to purchase if the sides of a bulletin bourd measures 3 yards by 2 yards

Answer by jorel555(1290)

If the board measures 3 x 2 yards, then the perimeter is:
(2x3)+(2x2)=6+4=10 yds. total. ☺☺☺☺